Please help me to solve this problem. Thank you.

thst Xmnd Y Lal XY be randem variobles; preve indepndent f) = giag,4 Where ake b, cxydihIR
Answer #1

TOPIC:Independence of random variables

Define the basic de dinition dons independence two sandom variables x and Y between as - It x and y are independent; then, EIF PART- Given that, 8x, y (x, y) = 9, (x). J₂ (2) ; acxeb, CRYLd. we need to prove that, X and Y are independent : ie. P (xL=) Now, we have P (x Ex, YEY E der Ctr, tu) dt, dt g, (+1). Iz (+2) . dt, dt ₂1 11:2.00m) { f29333 = P(x2x), 8 (427) ; & xey.7 ONLY IF PART :- Criven that x and r are independent We need to prove that x,y (267) = g(x). 9, 17) ; x 79. Now, since x and